# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

### 12 Mar 2024 v1

@@ 1 Introduction
Figure 1: One example of KG-LLM framework can be visually demonstrated with a simple task consisting of three stages. The original KG dataset is loaded in the first stage. In the second stage, we apply the graph preprocessing function to a randomly selected path in the graph, enabling the path to be represented as a chain-of-thought prompt. The third stage involves instruction fine-tuning the model to predict whether the first node is connected to the last node in the sub-KG.
@@
Despite these successes, our research highlights three major challenges that prior methodologies have not fully addressed, which our approach aims to solve. First, the predominant focus on discriminative, rather than generative, models and outcomes over reasoning processes underscores a gap in the existing methodologies, highlighting the need for models adept at leveraging reasoning to address multi-hop link prediction in KGs. Secondly, existing approaches predominantly focus on predicting links between two immediate nodes, leaving the field of multi-hop link prediction largely unexplored. This limitation affects the models’ ability to navigate and infer connections across extended node sequences. Lastly, the traditional models generally lack emergent abilities, making them less effective when faced with unseen tasks.
To bridge these gaps, we propose the Knowledge Graph Large Language Model Framework (KG-LLM Framework), a novel approach to multi-hop link prediction. As illustrated in Figure 1, envision a KG wherein nodes are interconnected through specific relations. Initially, our framework takes input of the original KG dataset. After preprocessing, a path in the KG (represented by the red links) is transformed into a chain-of-thought (CoT) prompt [27], enclosing through a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of a KG’s CoT prompt, determined by its path length and the number of nodes and connections, directly correlates with the intricacy of the multi-hop problem it represents. Through instruction fine-tuning (IFT) [25] of three Large Language Models (LLMs)—Flan-T5 [25], LlaMa2 [19], and Gemma [18]—our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) [29], the model not only improves but also acquires the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
@@ 3.3 KG-LLM Framework
Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL), as depicted in Figure 3. The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ emergent ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the standard and KG-LLM frameworks.
Figure 3: This figure illustrates our experimental overflow. We conducted eight tests for each model–Flan-T5, LLaMa2, and Gemma.
@@ 4 Experiments
$\bullet$ Q4: Can the application of ICL bolster the models’ emergent ability in multi-hop relation prediction tasks?
@@ 4.1 Experimental Setup
Metrics for Multi-hop Link Prediction.
Metrics for Multi-hop Relation Prediction.
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
In this analysis, we explore the models’ ability to predict relations not encountered during their training phase, exclude the traditional approach as it lacks emergent ability. We maintain consistency with the testing dataset used in the multi-hop link prediction task to ensure comparability and fairness. The difference lies in the modifications made to the instruction and prompt question presented to the model. Detailed specifications of these instructions are provided in the Appendix.
@@
Answer to Q3: The findings suggest that the KG-LLM framework marginally enhances the models’ emergent abilities. However, it would be premature to assert that our framework equips models with the ability to navigate unseen prompts. This could be attributed to the complexity and difficulty of the new instructions and options. With options no longer limited to a binary yes or no answer, the model may struggle to comprehend the updated instruction and effectively utilize the provided options.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
In addition to evaluating the performance of models on the relation prediction task without In-Context Learning (ICL), this study further explores the impact of incorporating ICL into the task. The ICL example is shown in the appendix. The comparative results, as illustrated in Table 5, reveal a significant improvement in the emergent abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4. In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.

### 13 Mar 2024 v2

@@ 1 Introduction
Figure 1: One example of KG-LLM framework can be visually demonstrated with a simple task consisting of two stages. In the first stage, we apply the graph preprocessing function to a randomly selected path in the knowledge graph, enabling the path to be represented as a chain-of-thought prompt. The second stage use instruction finetuning to finetune the LLM to predict whether the first node is connected to the last node in the sub-KG.
@@
Despite these successes, our research highlights three major challenges that prior methodologies have not fully addressed, which our approach aims to solve. First, the predominant focus on discriminative, rather than generative, models and outcomes over reasoning processes underscores a gap in the existing methodologies, highlighting the need for models adept at leveraging reasoning to address multi-hop link prediction in KGs. Secondly, existing approaches predominantly focus on predicting links between two immediate nodes, leaving the field of multi-hop link prediction largely unexplored. This limitation affects the models’ ability to navigate and infer connections across extended node sequences. Lastly, the traditional models generally lack generalization abilities, making them less effective when faced with unseen tasks.
To bridge these gaps, we propose the Knowledge Graph Large Language Model Framework (KG-LLM Framework), a novel approach to multi-hop link prediction. As illustrated in Figure 1, envision a KG wherein nodes are interconnected through specific relations. Initially, our framework takes input of the original KG dataset. After preprocessing, a path in the KG is transformed into a chain-of-thought (CoT) prompt [27], enclosing through a series of relational statements, which can be represented as {Node 1 (has relation_$x$ with) Node 2, Node 2 (has relation_$y$ with) Node 3, etc.}. The complexity of a KG’s CoT prompt, determined by its path length and the number of nodes and connections, directly correlates with the intricacy of the multi-hop problem it represents. Through instruction fine-tuning (IFT) [25] of three Large Language Models (LLMs)—Flan-T5 [25], LlaMa2 [19], and Gemma [18]—our framework is ready to enhance multi-hop link prediction performance during the testing phase. Moreover, by integrating in-context learning (ICL) [29], the model not only improves but also acquires the capacity to tackle unseen prompts, showcasing our method’s innovativeness in addressing multi-hop link prediction challenges.
@@ 3.3 KG-LLM Framework
Upon the completion of the training phases, we subject each model to two inference task sets, each comprising two sub-tests: non-In-Context Learning (non-ICL) and In-Context Learning (ICL). The primary set of inference tasks is centered around multi-hop link prediction. Conversely, the secondary set probes the models’ generalization ability in multi-hop relation prediction, particularly with previously unseen prompts. Through pre- and post-ICL evaluation within each task set, we aim to evaluate the impact of ICL integration across both the standard and KG-LLM frameworks.
@@ 4 Experiments
$\bullet$ Q4: Can the application of ICL bolster the models’ generalization ability in multi-hop relation prediction tasks?
@@ 4.1 Experimental Setup
Metrics for Multi-hop Link Prediction.
Metrics for Multi-hop Relation Prediction.
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
In this analysis, we explore the models’ ability to predict relations not encountered during their training phase, exclude the traditional approach as it lacks generalization ability. We maintain consistency with the testing dataset used in the multi-hop link prediction task to ensure comparability and fairness. The difference lies in the modifications made to the instruction and prompt question presented to the model. Detailed specifications of these instructions are provided in the Appendix.
@@
Answer to Q3: The findings suggest that the KG-LLM framework marginally enhances the models’ generalization abilities. However, it would be premature to assert that our framework equips models with the ability to navigate unseen prompts. This could be attributed to the complexity and difficulty of the new instructions and options. With options no longer limited to a binary yes or no answer, the model may struggle to comprehend the updated instruction and effectively utilize the provided options.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
In addition to evaluating the performance of models on the relation prediction task without In-Context Learning (ICL), this study further explores the impact of incorporating ICL into the task. The ICL example is shown in the appendix. The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the standard framework and the KG-LLM framework, in contrast to the findings presented in Table 4. In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.